### v1

# Sample Complexity of Composite Quantum Hypothesis Testing

Abstract
This paper investigates symmetric composite binary quantum hypothesis testing (QHT), where the goal is to determine which of two uncertainty sets contains an unknown quantum state. While asymptotic error exponents for this problem are well-studied, the finite-sample regime remains poorly understood. We bridge this gap by characterizing the sample complexity – the minimum number of state copies required to achieve a target error level. Specifically, we derive lower bounds that generalize the sample complexity of simple QHT and introduce new upper bounds for various uncertainty sets, including of both finite and infinite cardinalities. Notably, our upper and lower bounds match up to universal constants, providing a tight characterization of the sample complexity. Finally, we extend our analysis to the differentially private setting, establishing the sample complexity for privacy-preserving composite QHT.
I Introduction
Hypothesis testing is a fundamental problem in statistical inference where the goal is to determine which category a test object belongs to. Quantum hypothesis testing (QHT) extends this paradigm to the quantum realm, merging quantum mechanics with mathematical statistics, to address fundamental tasks in quantum information processing and communication [4, 3]. In its simplest form, a simple binary QHT problem is a state discrimination problem : Given $n$ copies of an unknown quantum state, a distinguisher must design a two-outcome quantum measurement to identify which of the two known states, $\rho$ or $\sigma$, was prepared [13, 16]. The performance of such a test is typically evaluated through two lenses: symmetric QHT, which seeks to minimize the average probability of error, and asymmetric QHT, which minimizes the type-II error subject to a fixed constraint on the type-I error [8].
While simple QHT involves distinguishing between individual states, many practical scenarios involve composite binary QHT. In this setting, the goal is to determine which of two uncertainty sets of quantum states contains the unknown state. This generalizes the problem to more complex, real-world conditions, where the exact state may not be known precisely, due to experimental limitations or environmental noise.
Existing research in both simple and composite QHT has primarily focused on asymptotic error exponents, which characterize the exponential rate of error decay as the number of state copies $n$ approaches infinity. For instance, the quantum Chernoff exponent for symmetric simple binary QHT was established in [2, 20] and recently generalized to the composite setting in [10]. Similarly, for asymmetric QHT, the quantum Stein exponent has been studied for simple QHT in [14, 21] and has been generalized to the composite setting under structural assumptions on the uncertainty sets in [12, 6, 5, 9, 11]. These results provide fundamental limits in the limit of infinite resources but do not fully capture the requirements of practical, resource-constrained quantum systems.
Recently, an alternative line of research has emerged to study QHT in the non-asymptotic setting by analysing the sample complexity. This refers to the minimum number of quantum state copies $n^{\ast}(\delta)$ required to achieve a desired target error level $\delta$. For simple symmetric QHT between quantum states $\rho_{1}$ and $\rho_{2}$, recent work [8] has characterized this complexity as $n^{\ast}(\delta)=\Theta\Bigl(\frac{\ln\frac{1}{\delta}}{-\ln F(\rho_{1},\rho_{2})}\Bigr)$ where $F(\rho_{1},\rho_{2})$, denotes the (square root) Uhlmann fidelity between the quantum states. This analysis has since been extended to simple QHT under local differential privacy constraints [7, 19], where the unknown quantum state is affected by a differentially private quantum channel before reaching the tester.
Building on these recent developments, this paper provides the first comprehensive characterization of the sample complexity of symmetric composite QHT. Our key contributions are as follows: First, we derive new lower and upper bounds on the sample complexity for various classes of uncertainty sets, including both finite and infinite cardinalities. Specifically, for testing a singleton pure uncertainty set $\mathcal{D}_{1}=\{\left|\psi\right\rangle\!\left\langle\psi\right|\}$ against a composite uncertainty set $\mathcal{D}_{2}$, we show that the sample complexity scales as $n^{\ast}(\delta)=\Theta\Bigl(\frac{\ln(\frac{1}{\delta})}{-\ln\sup_{\rho_{2}\in\mathcal{D}_{2}}\left\langle\psi\right|\rho_{2}\left|\psi\right\rangle}\Bigr)$. We generalize this result to the case when both the sets are finite, showing that the sample complexity is governed by the maximum pairwise fidelity between the sets, $F_{\max}=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}F(\rho_{1},\rho_{2})$, as $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(|\mathcal{D}_{1}|\,|\mathcal{D}_{2}|/\delta)}{-\ln F_{\max}}\Bigr)$. We also extend this analysis to infinite cardinality uncertainty sets under specific fidelity constraints. Secondly, we extend our framework to the locally differentially private setting, establishing the first sample complexity bounds for private composite QHT.
II Problem Setting
II-A Notation
For any bounded operator $A$, we denote $\operatorname{Tr}(A)$ as the trace of $A$, with its trace norm defined as $\|A\|_{1}:=\operatorname{Tr}(\sqrt{A^{\dagger}A})$. Throughout this paper, we use $\rho$ and $\sigma$ to denote quantum states represented as density matrices, i.e., positive semi-definite, unit-trace, Hermitian matrices, acting on a Hilbert space of dimension $d$. Then, $F(\rho,\sigma):=\|\sqrt{\rho}\sqrt{\sigma}\|_{1}$ defines the (square root) Uhlmann fidelity, and $d_{B}(\rho,\sigma):=\sqrt{\,2\left(1-F(\rho,\sigma)\right)}$ defines the Bures distance between the two states. We use $\dim_{\mathbb{R}}(\cdot)$ to denote the real dimension of the smallest affine space generated by the set $\cdot$’. If the set is a vector space, then $\dim_{\mathbb{R}}(\cdot)$ coincides with its standard linear dimension, i.e., the number of linearly independent vectors whose real span is the vector space.
II-B Simple Quantum Hypothesis Testing
In simple binary QHT, a quantum system is prepared in one of two known states $\rho_{1}^{\otimes}$ or $\rho_{2}^{\otimes n}$. The distinguisher is given access to $n$ identical copies of the state without a priori knowledge of which was prepared. The objective of the distinguisher is to determine the identity of the state by designing a two-outcome positive operator-valued measurement (POVM) $\{\Pi,I-\Pi\}$, where the outcome associated with the operator $0\leq\Pi\leq I$ corresponds to guessing $\rho_{1}^{\otimes n}$ and the outcome associated with $I-\Pi$ corresponds to guessing $\rho_{2}^{\otimes n}$.
The probability of error is determined by the Born’s rule. Specifically, the probability of erroneously guessing $\rho_{1}$ when the prepared state is $\rho_{2}$ is $\operatorname{Tr}(\Pi\rho_{2}^{\otimes n})$, while the converse error is $\operatorname{Tr}((I-\Pi)\rho_{1}^{\otimes n})$. Assuming prior probabilities $p$ and (1p)(1-p), where $p\in(0,1)$, for preparing the unknown state in state $\rho_{1}$ and $\rho_{2}$ respectively, the expected error probability of simple, symmetric QHT for a given POVM $\{\Pi,I-\Pi\}$ is defined as
$P_{e}(\Pi,p,\rho_{1}^{\otimes n},\rho_{2}^{\otimes n}):=p\operatorname{Tr}((I-\Pi)\rho_{1}^{\otimes n})+(1-p)\operatorname{Tr}(\Pi\rho_{2}^{\otimes n}).$ (1)
The optimal performance is found by minimizing over all possible POVMs. This minimum expected error probability is achieved by the Holevo-Helstrom measurement [13, 16] as
$P_{e,\min}(p,\rho_{1}^{\otimes n},$ $\rho_{2}^{\otimes n}):=\inf_{0\preceq\Pi\preceq I}P_{e}(\Pi,p,\rho_{1}^{\otimes n},\rho_{2}^{\otimes n})$ (2)
$=\tfrac{1}{2}\left(1-\left\|p\rho_{1}^{\otimes n}-(1-p)\rho_{2}^{\otimes n}\right\|_{1}\right).$ (3)
We now define the sample complexity for simple QHT.
Definition 1 (Sample Complexity of Simple QHT)
The $\delta$-sample complexity $n^{\ast}(p,\rho_{1},\rho_{2},\delta)$ for distinguishing between states $\rho_{1}$ and $\rho_{2}$ refers to the smallest number of quantum state copies required to achieve a minimum expected error probability of at most $\delta\in(0,1)$, i.e.,
$n^{\ast}(p,\rho_{1},\rho_{2},\delta):=\inf\{n\in\mathbb{N}:P_{e,\min}(p,\rho_{1}^{\otimes n},\rho_{2}^{\otimes n})\leq\delta\}.$
For singleton quantum states $\rho_{1}$ and $\rho_{2}$, recent work [8] establishes the following bounds on the sample complexity.
Theorem 2 ([8])
The following bounds hold for the sample complexity of simple QHT:
$\max\biggl\{\frac{\ln(p(1-p)/\delta)}{-2\ln F(\rho_{1},\rho_{2})},$ $\frac{1-\frac{\delta(1-\delta)}{p(1-p)}}{d_{B}(\rho_{1},\rho_{2})^{2}}\biggr\}\leq n^{\ast}(p,\rho_{1},\rho_{2},\delta)$
$\leq\biggl\lceil\inf_{s\in[0,1]}\frac{\ln\Bigl(\frac{p^{s}(1-p)^{1-s}}{\delta}\Bigr)}{-\ln\operatorname{Tr}(\rho_{1}^{s}\rho_{2}^{1-s})}\biggr\rceil.$ (4)
We next generalize this setting to composite hypotheses sets.
II-C Composite Quantum Hypothesis Testing
In composite QHT, the distinguisher tests if the unknown quantum state belongs to either of the two known uncertainty sets of quantum states: $\mathcal{D}_{1,n}$ or $\mathcal{D}_{2,n}$. We define these sets as
$\mathcal{D}_{i,n}:=\{\rho_{i}^{\otimes n}:\rho_{i}\in\mathcal{D}_{i}\},\quad\mbox{for}\hskip 2.84544pti=1,2,$ (5)
consisting of $n$-independent copies of quantum states drawn from compact sets $\mathcal{D}_{1}$ or $\mathcal{D}_{2}$ of density operators in a $d$-dimensional Hilbert space [5]. While simple QHT distinguishes between two fixed states, the composite framework requires discriminating between two sets of quantum states; it reduces to the simple case only when $\mathcal{D}_{i}$ are of unit cardinality.
In composite QHT, the distinguisher aims to design a two-outcome POVM $\{\Pi,I-\Pi\}$ that minimizes the probability of error, where the operator $\Pi$ corresponds to the hypothesis that the unknown state belongs to $\mathcal{D}_{1,n}$ and $I-\Pi$ to $\mathcal{D}_{2,n}$. However, unlike the simple QHT, the exact state within the uncertainty sets in which the quantum system is prepared is unknown, necessitating a worst-case approach. Therefore, we define the composite expected error probability of a given POVM as the supremum of the error probability over all possible state pairs within the uncertainty sets: [10]
$P_{e}(\Pi,p,\mathcal{D}_{1,n},\mathcal{D}_{2,n}):=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}P_{e}(\Pi,p,\rho_{1}^{\otimes n},\rho_{2}^{\otimes n}),$ (6)
where $P_{e}(\Pi,p,\rho_{1}^{\otimes n},\rho_{2}^{\otimes n})$ is defined as in (1) and (p,1p)(p,1-p) are the prior probabilities for each uncertainty set. The min-max expected error probability is then obtained as
$P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n}):=\inf_{0\preceq\Pi\preceq I}P_{e}(\Pi,p,\mathcal{D}_{1,n},\mathcal{D}_{2,n}).$ (7)
We can now extend the definition of sample complexity to composite QHT as follows.
Definition 3 (Sample Complexity of Composite QHT)
The $\delta$-sample complexity $n^{\ast}(p,\mathcal{D}_{1},\mathcal{D}_{2},\delta)$ for distinguishing between sets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ is defined as
$n^{\ast}(p,\mathcal{D}_{1},\mathcal{D}_{2},\delta):=\inf\{n\in\mathbb{N}:P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})\leq\delta\}.$
In the following sections, we provide the first study of the sample complexity of composite QHT. This analysis unveils how the sample complexity depends on key problem parameters, including the “similarity" of the uncertainty sets, the number of elements they contain, and the dimensionality of the underlying Hilbert space. For notational convenience, we abbreviate $n^{\ast}(p,\mathcal{D}_{1},\mathcal{D}_{2},\delta)$ (or $n^{\ast}(p,\rho_{1},\rho_{2},\delta)$) as $n^{\ast}(\delta)$ whenever the prior and the sets (or states) are clear from the context. Finally, we exclude the following trivial cases from our analysis, which generalizes the regimes identified in [8, Remark 2]. A proof can be found in Appendix -A.
Remark 4
If $\mathcal{D}_{1}\perp\mathcal{D}_{2}$, or $\delta\in[\tfrac{1}{2},1]$, or there exists an $s\in[0,1]$ such that $\delta\geq p^{s}(1-p)^{1-s}$, then $n^{\ast}(\delta)=1$. Also, if $\mathcal{D}_{1}\cap\mathcal{D}_{2}\neq\emptyset$ and $\delta<\min\{p,1-p\}$, then $n^{\ast}(\delta)=+\infty$.
III Main Results: Sample Complexity Bounds
In this section, we present our results, establishing upper and lower bounds on the sample complexity of composite QHT. We start by analysing the error probability, and use the insights as building blocks to derive the main bounds.
III-A Analysis of Min-max Expected Error Probability
We first provide an alternative characterization of the min-max expected error probability (7). To this end, we define the convex hulls of the uncertainty sets $\mathcal{D}_{i,n}$ in (5) as
$\mathcal{C}_{i,n}:=\operatorname{conv}(\mathcal{D}_{i,n})=\left\{\int\rho_{i}^{\otimes n}d\mu_{i}(\rho_{i}):\mu_{i}\in\mathcal{P}(\mathcal{D}_{i})\right\},$ (8)
where $\mathcal{P}(\mathcal{D}_{i})$ is the set of all probability measures on $\mathcal{D}_{i}$, for $i=1,2$. We then have the following result.
Proposition 5
The following relationship holds:
$P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})=\\ \sup_{\sigma_{i,n}\in\mathcal{C}_{i,n},i=1,2}\left(\tfrac{1}{2}-\tfrac{1}{2}\|p\sigma_{1,n}-(1-p)\sigma_{2,n}\|_{1}\right),$ (9)
where $\mathcal{C}_{i,n}$ is defined as in (8).
Proof:
Starting from the definition in (7), the following series of relationships hold:
$P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})$ $=\inf_{0\preceq\Pi\preceq I}\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n}}P_{e}(\Pi,p,\sigma_{1,n},\sigma_{2,n})$
$=\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n}}\inf_{0\preceq\Pi\preceq I}P_{e}(\Pi,p,\sigma_{1,n},\sigma_{2,n})$
$=\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n}}P_{e,\min}(p,\sigma_{1,n},\sigma_{2,n}),$
where in the first equality, we used linearity of $P_{e}$ in its argument quantum states, whereby the supremum is achieved at the extreme points of the convex hull. The second equality follows from applying Sion’s minimax theorem [5, Lemma A.1] to swap the order of $\sup$ and $\inf$, since $P_{e}(\Pi,p,\sigma_{1,n},\sigma_{2,n})$ is linear in $\Pi$, and the $\mathcal{C}_{i,n}$ are closed, compact and convex sets. The result then follows from the last equality via (3). ∎
Importantly, the following lemma shows that the error probability in Proposition˜5 decreases in the number of copies.
Lemma 6
Let $\Delta_{n}:=\inf_{\sigma_{i,n}\in\mathcal{C}_{i,n}}\|p\sigma_{1,n}-(1-p)\sigma_{2,n}\|_{1}$. Then, for all $n\in\mathbb{N}$, we have $\Delta_{n}\leq\Delta_{n+1}$ and consequently, $P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})\geq P_{e,\min}(p,\mathcal{D}_{1,n+1},\mathcal{D}_{2,n+1})$.
Proof:
The proof follows by using the partial trace over the $(n+1)$-th qubit, $\operatorname{Tr}_{n+1}(\cdot)$, which is a completely positive trace-preserving (CPTP) map, and noting that $\operatorname{Tr}_{n+1}(\mathcal{C}_{i,n+1})=\mathcal{C}_{i,n}.$ Then, for minimizers $\sigma_{i,n+1}^{*}\in\mathcal{C}_{i,n+1}$, $i=1,2$, of $\Delta_{n+1}$, we get
$\Delta_{n}$ $\leq\|\operatorname{Tr}_{n+1}(p\sigma_{1,n+1}^{*}-(1-p)\sigma_{2,n+1}^{*})\|_{1}$
$\leq\|p\sigma_{1,n+1}^{*}-(1-p)\sigma_{2,n+1}^{*}\|_{1}$
$=\inf_{\sigma_{i,n+1}\in\mathcal{C}_{i,n+1}}\|p\sigma_{1,n+1}-(1-p)\sigma_{2,n+1}\|_{1}=\Delta_{n+1},$
where the first inequality follows because $\operatorname{Tr}_{n+1}(\sigma_{i,n+1}^{*})\in\mathcal{C}_{i,n}$ and the second inequality uses that $\|\operatorname{Tr}_{n+1}(A)\|_{1}\leq\|A\|_{1}$ for a CPTP map acting on a Hermitian operator $A$. ∎
We next present a series of upper bounds on the min-max expected error probability in (9), which will be used later.
Proposition 7
The following series of upper bounds hold:
$P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})$
$\leq\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n},i=1,2}\inf_{0\leq s\leq 1}p^{s}(1-p)^{1-s}\operatorname{Tr}(\sigma_{1,n}^{s}\sigma_{2,n}^{1-s})$ (10)
$\leq\sqrt{p(1-p)}\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n},i=1,2}\operatorname{Tr}(\sqrt{\sigma_{1,n}}\sqrt{\sigma_{2,n}})$ (11)
$\leq\sqrt{p(1-p)}\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n},i=1,2}F(\sigma_{1,n},\sigma_{2,n}).$ (12)
Proof:
The first inequality uses the well-known quantum Chernoff Bound [1, Theorem 1], while the second inequality follows by choosing $s=1/2$. The last inequality follows from [2, Eq. 28] and the definition of Uhlmann fidelity. ∎
III-B Lower Bounds on Sample Complexity
We now present lower bounds on the $\delta$-sample complexity of composite QHT, generalizing the results in [8]. We define
$F_{\max}:=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}F(\rho_{1},\rho_{2})$ (13)
as the maximum fidelity between pairs of quantum states within the uncertainty sets.
Theorem 8
Let $\mathcal{D}_{1},\mathcal{D}_{2}$ be arbitrary compact sets of quantum states. Then, for $p,\delta\in(0,1)$,
$\max\left\{\frac{\ln\left(\frac{p(1-p)}{\delta}\right)}{-2\ln F_{\max}},\\ \frac{1-\frac{\delta(1-\delta)}{p(1-p)}}{\inf_{\rho_{i}\in\mathcal{D}_{i}}d_{B}^{2}(\rho_{1},\rho_{2})}\right\}\leq n^{\ast}(\delta).$ (14)
Proof:
Since the min-max error probability is defined in a worst-case sense over the uncertainty sets, we have $n^{\ast}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq n^{\ast}(\delta,p,\rho_{1},\rho_{2})$ for any pair $(\rho_{1},\rho_{2})$, and hence $n^{\ast}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq\sup_{\rho_{i}\in\mathcal{D}_{i}}n^{\ast}(\delta,p,\rho_{1},\rho_{2})$. Therefore, the desired result follows from the lower bound in Theorem˜2 by taking the supremum over all pairs of states $(\rho_{1},\rho_{2})$. ∎
Theorem˜8 shows that the lower bound of the sample complexity depends on the least distinguishable pair of states from $\mathcal{D}_{1}\times\mathcal{D}_{2}$. For singleton sets, Theorem˜8 collapses to the lower bound in Theorem˜2.
III-C Upper Bounds on Sample Complexity
We now present upper bounds on the sample complexity by considering uncertainty sets of increasing “complexity”.
III-C1 Quantum State Verification Problem
Consider the task of quantum state verification (QSV) [22], where $\mathcal{D}_{1}$ consists of a singleton pure state $\left|\psi\right\rangle\!\left\langle\psi\right|$, and $\mathcal{D}_{2}$ be any compact set of quantum states satisfying $\left|\psi\right\rangle\!\left\langle\psi\right|\notin\mathcal{D}_{2}$. The following theorem provides an upper bound on the sample complexity of QSV.
Theorem 9
Let $\mathcal{D}_{1}=\{\left|\psi\right\rangle\!\left\langle\psi\right|\}$ be a singleton pure state and $\mathcal{D}_{2}$ be a compact set of quantum states satisfying $\mathcal{D}_{1}\cap\mathcal{D}_{2}=\emptyset$. Then, for $p,\delta\in(0,1)$,
$n^{\ast}(\delta)\leq\left\lceil\frac{\ln\Bigl(\frac{1-p}{\delta}\Bigr)}{-\ln\sup_{\rho_{2}\in\mathcal{D}_{2}}\left\langle\psi\right|\rho_{2}\left|\psi\right\rangle}\right\rceil.$ (15)
For fixed $p$, (14) and (15) together imply that the sample complexity for QSV scales as $n^{\ast}(\delta)=\Theta\Bigl(\frac{\ln(1/\delta)}{-\ln\sup_{\rho_{2}\in\mathcal{D}_{2}}\left\langle\psi\right|\rho_{2}\left|\psi\right\rangle}\Bigr).$
Proof:
The proof of (15) follows from using the upper bound in (10) by noting that $\sigma_{1,n}^{s}=(\left|\psi\right\rangle\!\left\langle\psi\right|^{\otimes n})^{s}=\left|\psi\right\rangle\!\left\langle\psi\right|^{\otimes n}$ for $s\in[0,1]$. Choosing $s=0$, then yields that
$P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})$ $\leq(1-p)\sup_{\sigma_{2,n}\in\mathcal{C}_{2,n}}\operatorname{Tr}(\sigma_{2,n}\left|\psi\right\rangle\!\left\langle\psi\right|^{\otimes n})$
$=(1-p)\sup_{\rho_{2}\in\mathcal{D}_{2}}\operatorname{Tr}(\rho_{2}\left|\psi\right\rangle\!\left\langle\psi\right|)^{n},$
where the equality follows since the supremum is attained at an extreme point of the convex hull. The choice of $n$ in (15) satisfies $P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})\leq\delta$, completing the proof. ∎
The theorem above shows that the sample complexity for QSV problem scales with the maximum overlap between the pure state $\left|\psi\right\rangle\!\left\langle\psi\right|$ and the states in the uncertainty set $\mathcal{D}_{2}$. When $\mathcal{D}_{2}$ is also singleton, the bound (15) recovers the sample complexity upper bound of simple QHT in (4) with $s=0$ and $\rho=\left|\psi\right\rangle\!\left\langle\psi\right|$. We note that the choice of $s=0$ is in fact the minimizer of (10) when $p\geq\frac{1}{2}$ (see Appendix -B for a proof). Lastly, we note that the error exponent of QSV problem evaluates to
$\hskip-8.5359pt\lim_{n\rightarrow\infty}\frac{-\ln P_{e,\min}(p,\{\left|\psi\right\rangle\!\left\langle\psi\right|\},\mathcal{D}_{2,n})}{n}=-\!\ln\!\sup_{\rho_{2}\in\mathcal{D}_{2}}\left\langle\psi\right|\rho_{2}\left|\psi\right\rangle,$
recovering the composite quantum Chernoff exponent for $s=0$ [10]. This is consistent with the known singleton pure–mixed case, where the optimal exponent is attained at $s=0$ [1].
III-C2 Uncertainty Sets of Finite Cardinality
We now extend singleton uncertainty sets to uncertainty sets of finite cardinality, i.e., $|\mathcal{D}_{i}|=m_{i}<\infty$ for $i=1,2$. To upper bound the sample complexity, we first derive the following result.
Lemma 10
Suppose that $|\mathcal{D}_{i}|=m_{i}<\infty$. Then,
$\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n},i=1,2}F(\sigma_{1,n},\sigma_{2,n})\leq\sqrt{m_{1}m_{2}}F_{\max}^{n}.$ (16)
Proof:
For $i=1,2$, let $\sigma_{i,n}\in\mathcal{C}_{i,n}$ denote quantum states in the respective convex hulls, which can be written equivalently as $\sigma_{1,n}=\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n}$ and $\sigma_{2,n}=\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n},$ where $\rho_{1,j}\in\mathcal{D}_{1}$, $\rho_{2,k}\in\mathcal{D}_{2}$ and $\{p_{j}\},\{q_{k}\}$ denote probability vectors. We then have
$F(\sigma_{1,n},\sigma_{2,n})$ $=\left\|\sqrt{\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n}}\sqrt{\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n}}\right\|_{1}$
$\leq\left\|\sum_{j=1}^{m_{1}}\sqrt{p_{j}\rho_{1,j}^{\otimes n}}\sum_{k=1}^{m_{2}}\sqrt{q_{k}\rho_{2,k}^{\otimes n}}\right\|_{1}$
$=\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
$\leq\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
$\leq\sqrt{m_{1}m_{2}}F_{\max}^{n},$
where the first inequality follows by using Rotfel’d trace inequality twice, since the square root is concave [17, Theorem 1.1] and the last inequality follows since $\sum_{j=1}^{m_{1}}\sqrt{p_{j}}\leq\sqrt{m_{1}}$. ∎
We now derive the following upper bound using Lemma˜10.
Theorem 11
Let $|\mathcal{D}_{i}|=m_{i}<\infty$ for $i=1,2$. Then, for every $p\in(0,1)$ and $\delta\in(0,1)$,
$n^{\ast}(\delta)\leq\left\lceil\frac{\ln\Bigl(\frac{\sqrt{m_{1}m_{2}p(1-p)}}{\delta}\Bigr)}{-\ln F_{\max}}\right\rceil.$ (17)
Furthermore, for fixed $p$, (14) and (17) together imply that $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(m_{1}m_{2}/\delta)}{-\ln F_{\max}}\Bigr).$
Proof:
The proof of (17) follows by combining the upper bound in (12) with Lemma˜10 and by verifying that the chosen value of $n$ satisfies $P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})\leq\delta$. ∎
III-C3 Uncertainty Sets of Infinite Cardinality
Lastly, we consider the most general setting where the uncertainty sets $\mathcal{D}_{i}$ are of infinite cardinality and derive upper bounds on the sample complexity. We first extend Lemma˜10 to this setting.
Lemma 12
Let $\mathcal{D}_{1},\mathcal{D}_{2}$ be compact sets of quantum states possibly of infinite cardinality. Then,
$\hskip-6.25963pt\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n}}F(\sigma_{1,n},\sigma_{2,n})\leq\left(\max_{i=1,2}\dim_{\mathbb{R}}(\mathcal{D}_{i,n})+1\right)F_{\max}^{n}.$ (18)
Moreover, we have for each $i=1,2$, $\dim_{\mathbb{R}}(\mathcal{D}_{i,n})+1\leq\binom{n+d^{2}-1}{n}$, where $d$ is the dimension of the Hilbert space.
Proof:
The proof uses Carathéodory’s theorem [23, Theorem 0.0.1] to write each element of $\mathcal{C}_{i,n}$ as a convex combination of $\dim_{\mathbb{R}}(\mathcal{C}_{i,n})+1$ elements of $\mathcal{D}_{i,n}$, then apply a version of Lemma˜10 with $\dim_{\mathbb{R}}(\mathcal{C}_{i,n})=\dim_{\mathbb{R}}(\mathcal{D}_{i,n})$. For details, see Appendix -C. ∎
Comparing Lemma˜10 and Lemma˜12, we note that the latter yields a tighter bound whenever $\sqrt{|\mathcal{D}_{1}|\,|\mathcal{D}_{2}|}>\max_{i=1,2}\dim_{\mathbb{R}}(\mathcal{D}_{i,n})+1$. However, Lemma˜12 shows that the multiplicative factor $\dim_{\mathbb{R}}(\mathcal{D}_{i,n})+1$ scales as $\operatorname{poly}(n)$. Consequently, the following upper bound on the probability of error obtained by combining the above lemma with (12),
$\hskip-5.69046ptP_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})\leq\sqrt{p(1-p)}\binom{n+d^{2}-1}{n}F_{\max}^{n},$ (19)
is not necessarily monotonically decreasing in $n$, contradicting Lemma˜6. To derive analytical sample complexity in the infinite cardinality setting, we restrict uncertainty sets to satisfy the max-fidelity constraint $F_{\max}\leq\frac{2}{d^{2}+1}$, which ensures that the upper bound in (19) decreases monotonically with $n$ (see Appendix -D for the proof). The following theorem presents an upper bound on the sample complexity for qubit uncertainty sets (i.e., $d=2$) satisfying the max-fidelity constraint.
Theorem 13
Let $\mathcal{D}_{1},\mathcal{D}_{2}$ be compact sets of qubit states (i.e., $d=2$) of infinite cardinality satisfying $F_{\max}\leq\frac{2}{5}$. For $p,\delta\in(0,1)$, we have
$n^{\ast}(\delta)\leq\left\lceil\inf_{a\in(0,\ln(5/2))}\frac{\ln\frac{\sqrt{p(1-p)}}{\delta}+\ln\frac{9e^{3(a-1)}}{2a^{3}}}{-\ln{F_{\max}}-a}\right\rceil.$ (20)
Additionally, for fixed $p$, (14) and (20) together imply that for uncertainty sets satisfying max-fidelity constraints, the sample complexity scales as $n^{\ast}(\delta)=\Theta\Bigl(\frac{\ln(\frac{1}{\delta})}{-\ln F_{\max}}\Bigr)$.
Proof:
For $d=2$, we can upper bound the combinatorial term in (19) as $\binom{n+3}{n}\leq\frac{1}{6}(n+3)^{3}$ and use the inequality $x^{k}\leq\left(\frac{k}{ae}\right)^{k}e^{ax}$ (proved in Appendix -E) with $x=n+3$, $k=3$ and $0<a<\ln{\frac{5}{2}}$, the latter ensuring the bound decays with $n$. This results in
$P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})\leq\frac{\sqrt{p(1-p)}}{6}\left(\frac{3}{ae}\right)^{3}e^{3a}\left(e^{a}F_{\max}\right)^{n}.$
Choosing $n$ as in (20) ensures $P_{e,\min}(p,\mathcal{D}_{1,n},\mathcal{D}_{2,n})\leq\delta$, completing the proof. ∎
IV Differentially Private Composite QHT
We now extend the previous results to the setting when the unknown quantum state is pre-processed by a noisy quantum channel before we receive it for testing. Specifically, we focus on the class of $\varepsilon$-locally differentially private quantum (LDPQ) channels, defined as follows.
Definition 14
A CPTP map $\mathcal{M}$ is $\varepsilon$-LDPQ for $\varepsilon\geq 0$ if
$\sup_{\rho_{1},\rho_{2}}E_{e^{\varepsilon}}\big(\mathcal{M}(\rho_{1})\|\mathcal{M}(\rho_{2})\big)=0,$ (21)
where $E_{\gamma}(\rho_{1},\rho_{2})=\operatorname{Tr}\left[(\rho_{1}-\gamma\rho_{2})_{+}\right]$ is the quantum hockey-stick divergence [15].
Let $\mathcal{M}$ be an $\varepsilon$-LDPQ channel applied independently to each of the $n$-copies of the unknown quantum state. We then denote the locally private uncertainty sets and their corresponding convex hulls respectively as
$\mathcal{D}_{i,n}^{\mathcal{M}}:=\{\mathcal{M}(\rho_{i})^{\otimes n}:\rho_{i}\in\mathcal{D}_{i}\},\quad\mathcal{C}_{i,n}^{\mathcal{M}}:=\operatorname{conv}(\mathcal{D}_{i,n}^{\mathcal{M}}),$
for $i=1,2$, and define the $(\varepsilon,\delta)$-LDPQ sample complexity
$n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2}):=\!\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}n^{\ast}(\delta,p,\mathcal{D}_{1}^{\mathcal{M}},\mathcal{D}_{2}^{\mathcal{M}}),$ (22)
where $\operatorname{LDP}_{\varepsilon}$ is the set of all $\varepsilon$-LDPQ channels. We abbreviate the above notation as $n^{*}_{\varepsilon}(\delta)$ when the prior and the uncertainty sets are clear.
The following result presents a lower bound on this sample complexity, where we define $H_{1/2}(\rho_{1}\|\rho_{2}):=2\left(1-\operatorname{Tr}[\sqrt{\rho_{1}}\sqrt{\rho_{2}}]\right)$ as the Hellinger divergence of order $1/2$.
Proposition 15
Let $\mathcal{D}_{1},\mathcal{D}_{2}$ be compact sets of quantum states such that $H_{1/2}(\rho_{1}\|\rho_{2})\leq 1$ for every $\rho_{i}\in\mathcal{D}_{i}$, $i=1,2$. Then for $p,\delta\in(0,1)$ and $\varepsilon\geq 0$, we have
$\max\biggl\{\frac{\left(1-\tfrac{\delta}{p}\right)^{2}}{e^{-\varepsilon}(e^{\varepsilon}-1)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i},i=1,2}E_{1}(\rho_{1}\|\rho_{2})^{2}},$
$\frac{(e^{\varepsilon}+1)\log\frac{p(1-p)}{\delta}}{2(e^{\varepsilon}-1)\inf_{\rho_{i}\in\mathcal{D}_{i},i=1,2}H_{1/2}(\rho_{1}\|\rho_{2})}\biggr\}\leq n^{\ast}_{\varepsilon}(\delta).$ (23)
Proof:
Using [7, Lemma 1], we have that $\ln F(\rho_{1},\rho_{2})\leq-\ln\left(1-\frac{1}{2}H_{1/2}(\rho_{1}\|\rho_{2})\right)\leq H_{1/2}(\rho_{1}\|\rho_{2})$, where we used $\ln(1-x)\leq 2x$ for all $x\in[0,\frac{1}{2}]$. Applying this to the lower bound of Theorem˜2, and using [7, (4.12)] gives $n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})\geq\frac{(e^{\varepsilon}+1)\ln\frac{p(1-p)}{\delta}}{2(e^{\varepsilon}-1)H_{1/2}(\rho_{1}\|\rho_{2})}$. Additionally, [7, (4.25)] yields that $n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})\geq\frac{(1-\delta/p)^{2}}{e^{-\varepsilon}(e^{\varepsilon}-1)^{2}E_{1}(\rho_{1}\|\rho_{2})^{2}}$. The proof then follows from noting that $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$ for any $(\rho_{1},\rho_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$, whereby we get $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq\sup_{\rho_{i}\in\mathcal{D}_{i}}n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$. ∎
The above result directly generalizes the sample complexity of differentially private simple QHT, recovering [7, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
However, as in the previous section, the real challenge is to derive tight upper bounds on the sample complexity. To this end, we derive the following result.
Lemma 16
For the setting of Proposition˜15, we have
$\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2}\leq-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\ln F^{\mathcal{M}}_{\max},$ (24)
where $F^{\mathcal{M}}_{\max}=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}F(\mathcal{M}(\rho_{1}),\mathcal{M}(\rho_{2}))$.
Proof:
Let $\mathcal{B}$ be the $\varepsilon$-LDP channel as constructed in the proof of [7, Theorem 4.2]. Then, we have
$-\ln\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$ (25)
$\geq 1-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$ (26)
$\geq\frac{1}{2}\sup_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\mathcal{M}(\rho_{1})\|\mathcal{M}(\rho_{2}))^{2}$ (27)
$\geq\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2},$ (28)
where in (26), we use that $1-x\leq-\ln x$ for all $x\in[0,1]$. (27) uses [7, Equation 4.7], and (28) uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$. ∎
Lemma˜16 establishes a lower bound on $\ln F_{\max}$ after $\varepsilon$-LDPQ pre-processing. Since all upper bounds derived in Section˜III are monotonically decreasing functions in $\ln F_{\max}$, combining Lemma˜16 with these expressions, yields valid upper bounds for the private setting. In particular, the following theorem elucidates this for finite cardinality sets.
Theorem 17
Let $\mathcal{D}_{1}$, $\mathcal{D}_{2}$ be compact sets of quantum states satisfying $|\mathcal{D}_{i}|=m_{i}<\infty$ for $i=1,2$. For $\varepsilon\geq 0$, and $p,\delta\in(0,1)$, we have
$n^{\ast}_{\varepsilon}(\delta)\leq\left\lceil\left(\frac{e^{\varepsilon}+1}{e^{\varepsilon}-1}\right)^{2}\frac{2\ln\Bigl(\tfrac{\sqrt{m_{1}m_{2}p(1-p)}}{\delta}\Bigr)}{\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2}}\right\rceil.$ (29)
Lastly, for fixed $p$ and $\delta$, (23) and (29) together imply that
$\Omega\left(\frac{1}{e^{-\varepsilon}(e^{\varepsilon}-1)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2}}\right)\leq n^{\ast}_{\varepsilon}(\delta)\\ \leq\mathcal{O}\left(\left(\frac{e^{\varepsilon}+1}{e^{\varepsilon}-1}\right)^{2}\frac{\ln(m_{1}m_{2})}{\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2}}\right).$ (30)
V Conclusion
This work characterized the sample complexity of binary composite QHT, establishing bounds that highlight the roles of set cardinality, maximum pairwise fidelity, and Hilbert space dimensionality. We extended these results to the locally differentially private regime, providing a robust framework for state discrimination under privacy constraints. Future research may generalize this to $M$-ary composite hypotheses and composite channel discrimination, or refine the analysis through an extension to the asymmetric setting.
Acknowledgments
STJ received funding from EPSRC Quantum Technologies Career Acceleration Fellowship (UKRI1218).

### v2

@@
 Specifically, for testing a singleton pure uncertainty set $\mathcal{D}_{1}=\{\left|\psi\right\rangle\!\left\langle\psi\right|\}$ against a composite uncertainty set $\mathcal{D}_{2}$, we show that the sample complexity scales as $n^{\ast}(\delta)=\Theta\Bigl(\frac{\ln(\frac{1}{\delta})}{-\ln\sup_{\rho_{2}\in\mathcal{D}_{2}}\left\langle\psi\right|\rho_{2}\left|\psi\right\rangle}\Bigr)$.
-We generalize this result to the case when both the sets are finite, showing that the sample complexity is governed by the maximum pairwise fidelity between the sets, $F_{\max}=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}F(\rho_{1},\rho_{2})$, as $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(|\mathcal{D}_{1}|\,|\mathcal{D}_{2}|/\delta)}{-\ln F_{\max}}\Bigr)$.
+We generalize this result to the case when both the sets are finite, showing that the sample complexity is governed by the maximum pairwise fidelity between the sets, $F_{\max}=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}F(\rho_{1},\rho_{2})$, as $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(\sqrt{|\mathcal{D}_{1}|\,|\mathcal{D}_{2}|/}\delta)}{-\ln F_{\max}}\Bigr)$.
 We also extend this analysis to infinite cardinality uncertainty sets under specific fidelity constraints.
@@
 II-B Simple Quantum Hypothesis Testing
-In simple binary QHT, a quantum system is prepared in one of two known states $\rho_{1}^{\otimes}$ or $\rho_{2}^{\otimes n}$.
+In simple binary QHT, a quantum system is prepared in one of two known states $\rho_{1}^{\otimes n}$ or $\rho_{2}^{\otimes n}$.
 The distinguisher is given access to $n$ identical copies of the state without a priori knowledge of which was prepared.
@@
 Lemma 10
-Suppose that $|\mathcal{D}_{i}|=m_{i}<\infty$. Then,
+Let $|\mathcal{D}_{i}|=m_{i}<\infty$ for $i=1,2$. Then,
 $\sup_{\sigma_{i,n}\in\mathcal{C}_{i,n},i=1,2}F(\sigma_{1,n},\sigma_{2,n})\leq\sqrt{m_{1}m_{2}}F_{\max}^{n}.$ (16)
@@
 $\leq\left\|\sum_{j=1}^{m_{1}}\sqrt{p_{j}\rho_{1,j}^{\otimes n}}\sum_{k=1}^{m_{2}}\sqrt{q_{k}\rho_{2,k}^{\otimes n}}\right\|_{1}$
-$=\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
-$\leq\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
+$\leq\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
+$=\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
 $\leq\sqrt{m_{1}m_{2}}F_{\max}^{n},$
@@
 $n^{\ast}(\delta)\leq\left\lceil\frac{\ln\Bigl(\frac{\sqrt{m_{1}m_{2}p(1-p)}}{\delta}\Bigr)}{-\ln F_{\max}}\right\rceil.$ (17)
-Furthermore, for fixed $p$, (14) and (17) together imply that $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(m_{1}m_{2}/\delta)}{-\ln F_{\max}}\Bigr).$
+Furthermore, for fixed $p$, (14) and (17) together imply that $\Omega\Bigl(\frac{\ln(1/\delta)}{-\ln F_{\max}}\Bigr)\leq n^{\ast}(\delta)\leq\mathcal{O}\Bigl(\frac{\ln(\sqrt{m_{1}m_{2}}/\delta)}{-\ln F_{\max}}\Bigr).$
 Proof:
@@
 where $E_{\gamma}(\rho_{1},\rho_{2})=\operatorname{Tr}\left[(\rho_{1}-\gamma\rho_{2})_{+}\right]$ is the quantum hockey-stick divergence [15].
-Let $\mathcal{M}$ be an $\varepsilon$-LDPQ channel applied independently to each of the $n$-copies of the unknown quantum state.
-We then denote the locally private uncertainty sets and their corresponding convex hulls respectively as
-$\mathcal{D}_{i,n}^{\mathcal{M}}:=\{\mathcal{M}(\rho_{i})^{\otimes n}:\rho_{i}\in\mathcal{D}_{i}\},\quad\mathcal{C}_{i,n}^{\mathcal{M}}:=\operatorname{conv}(\mathcal{D}_{i,n}^{\mathcal{M}}),$
+Let $\mathcal{M}$ be an $\varepsilon$-LDPQ channel applied independently to each of the $n$-copies of the unknown quantum state. We then denote the locally private uncertainty sets as
+$\mathcal{D}_{i,n}^{\mathcal{M}}:=\{\mathcal{M}(\rho_{i})^{\otimes n}:\rho_{i}\in\mathcal{D}_{i}\},$
 for $i=1,2$, and define the $(\varepsilon,\delta)$-LDPQ sample complexity
@@
 Using [7, Lemma 1], we have that $\ln F(\rho_{1},\rho_{2})\leq-\ln\left(1-\frac{1}{2}H_{1/2}(\rho_{1}\|\rho_{2})\right)\leq H_{1/2}(\rho_{1}\|\rho_{2})$, where we used $\ln(1-x)\leq 2x$ for all $x\in[0,\frac{1}{2}]$.
-Applying this to the lower bound of Theorem˜2, and using [7, (4.12)] gives $n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})\geq\frac{(e^{\varepsilon}+1)\ln\frac{p(1-p)}{\delta}}{2(e^{\varepsilon}-1)H_{1/2}(\rho_{1}\|\rho_{2})}$.
-Additionally, [7, (4.25)] yields that $n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})\geq\frac{(1-\delta/p)^{2}}{e^{-\varepsilon}(e^{\varepsilon}-1)^{2}E_{1}(\rho_{1}\|\rho_{2})^{2}}$.
+Applying this to the lower bound of Theorem˜2, and using [7, (4.12)] gives n^∗_ε(δ,p,ρ_1,ρ_2) ≥(eε+1)lnp(1-p)δ2(eε-1)H1/2(ρ1ρ2).
+Additionally, [7, (4.25)] yields that n^∗_ε(δ,p,ρ_1,ρ_2) ≥(1-δ/p)2e-ε(eε-1)2E1(ρ1ρ2)2.
 The proof then follows from noting that $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$ for any $(\rho_{1},\rho_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$, whereby we get $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq\sup_{\rho_{i}\in\mathcal{D}_{i}}n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$.
 ∎
-The above result directly generalizes the sample complexity of differentially private simple QHT, recovering [7, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
+The above result directly generalizes the sample complexity of locally differentially private simple QHT, recovering [7, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
 However, as in the previous section, the real challenge is to derive tight upper bounds on the sample complexity. To this end, we derive the following result.
 Lemma 16
-For the setting of Proposition˜15, we have
-$\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2}\leq-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\ln F^{\mathcal{M}}_{\max},$ (24)
+Consider the setting of Proposition˜15. Then, the following inequality holds,
+$\frac{1}{2}\biggl(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\biggr)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2}\leq-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\ln F^{\mathcal{M}}_{\max},$ (24)
 where $F^{\mathcal{M}}_{\max}=\sup_{\rho_{i}\in\mathcal{D}_{i},i=1,2}F(\mathcal{M}(\rho_{1}),\mathcal{M}(\rho_{2}))$.
@@
 Let $\mathcal{B}$ be the $\varepsilon$-LDP channel as constructed in the proof of [7, Theorem 4.2]. Then, we have
-$-\ln\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$ (25)
-$\geq 1-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$ (26)
-$\geq\frac{1}{2}\sup_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\mathcal{M}(\rho_{1})\|\mathcal{M}(\rho_{2}))^{2}$ (27)
-$\geq\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2},$ (28)
-where in (26), we use that $1-x\leq-\ln x$ for all $x\in[0,1]$.
-(27) uses [7, Equation 4.7], and (28) uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$.
+$-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\ln F^{\mathcal{M}}_{\max}$ $\geq 1-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$
+$\geq\frac{1}{2}\sup_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\mathcal{M}(\rho_{1})\|\mathcal{M}(\rho_{2}))^{2}$
+$\geq\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2},$
+where the first inequality uses $1-x\leq-\ln x$ for all $x\in[0,1]$, the second uses [7, Equation 4.7], and the last inequality uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$.
 ∎

### v3

@@
 For notational convenience, we abbreviate $n^{\ast}(p,\mathcal{D}_{1},\mathcal{D}_{2},\delta)$ (or $n^{\ast}(p,\rho_{1},\rho_{2},\delta)$) as $n^{\ast}(\delta)$ whenever the prior and the sets (or states) are clear from the context.
-Finally, we exclude the following trivial cases from our analysis, which generalizes the regimes identified in [8, Remark 2].
+Finally, we exclude the following trivial cases from our analysis, which generalizes the regimes identified in [9, Remark 2].
 A proof can be found in Appendix -A.
@@
 where in the first equality, we used linearity of $P_{e}$ in its argument quantum states, whereby the supremum is achieved at the extreme points of the convex hull.
-The second equality follows from applying Sion’s minimax theorem [5, Lemma A.1] to swap the order of $\sup$ and $\inf$, since $P_{e}(\Pi,p,\sigma_{1,n},\sigma_{2,n})$ is linear in $\Pi$, and the $\mathcal{C}_{i,n}$ are closed, compact and convex sets.
+The second equality follows from applying Sion’s minimax theorem [6, Lemma A.1] to swap the order of $\sup$ and $\inf$, since $P_{e}(\Pi,p,\sigma_{1,n},\sigma_{2,n})$ is linear in $\Pi$, and the $\mathcal{C}_{i,n}$ are closed, compact and convex sets.
 The result then follows from the last equality via (3).
@@
 The first inequality uses the well-known quantum Chernoff Bound [1, Theorem 1], while the second inequality follows by choosing $s=1/2$.
-The last inequality follows from [2, Eq.
+The last inequality follows from [3, Eq.
 28] and the definition of Uhlmann fidelity.
@@
 Proof:
-For $i=1,2$, let $\sigma_{i,n}\in\mathcal{C}_{i,n}$ denote quantum states in the respective convex hulls, which can be written equivalently as $\sigma_{1,n}=\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n}$ and $\sigma_{2,n}=\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n},$ where $\rho_{1,j}\in\mathcal{D}_{1}$, $\rho_{2,k}\in\mathcal{D}_{2}$ and $\{p_{j}\},\{q_{k}\}$ denote probability vectors.
-We then have
-$F(\sigma_{1,n},\sigma_{2,n})$ $=\left\|\sqrt{\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n}}\sqrt{\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n}}\right\|_{1}$
-$\leq\left\|\sum_{j=1}^{m_{1}}\sqrt{p_{j}\rho_{1,j}^{\otimes n}}\sum_{k=1}^{m_{2}}\sqrt{q_{k}\rho_{2,k}^{\otimes n}}\right\|_{1}$
-$\leq\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
-$=\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
+For $i=1,2$, let $\sigma_{i,n}\in\mathcal{C}_{i,n}$ denote any quantum states in the respective convex hulls, which can be written equivalently as $\sigma_{1,n}=\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n}$ and $\sigma_{2,n}=\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n},$ where $\rho_{1,j}\in\mathcal{D}_{1}$, $\rho_{2,k}\in\mathcal{D}_{2}$ and $\{p_{j}\},\{q_{k}\}$ denote probability vectors.
+For any $\sigma_{i,n}\in\mathcal{C}_{i,n}$, the following set of relations hold,
+$F(\sigma_{1,n},\sigma_{2,n})$ $=F\left(\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n},\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n}\right)$
+$\leq\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}F(p_{j}\rho_{1,j}^{\otimes n},q_{k}\rho_{2,k}^{\otimes n})$
+$=\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}F(\rho_{1,j},\rho_{2,k})^{n}$
+$\leq\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}F(\rho_{1},\rho_{2})^{n}$
 $\leq\sqrt{m_{1}m_{2}}F_{\max}^{n},$
-where the first inequality follows by using Rotfel’d trace inequality twice, since the square root is concave [17, Theorem 1.1] and the last inequality follows since $\sum_{j=1}^{m_{1}}\sqrt{p_{j}}\leq\sqrt{m_{1}}$.
+where the first inequality follows by sub-additivity of Fidelity [2, Lemma 4.9], and the last inequality follows by Cauchy-Schwarz: $\sum_{j=1}^{m_{1}}(\sqrt{p_{j}}\cdot 1)\leq\sqrt{(\sum_{j=1}^{m_{1}}p_{j})(\sum_{j=1}^{m_{1}}1^{2})}=\sqrt{m_{1}}$.
 ∎
@@
 Proof:
-Using [7, Lemma 1], we have that $\ln F(\rho_{1},\rho_{2})\leq-\ln\left(1-\frac{1}{2}H_{1/2}(\rho_{1}\|\rho_{2})\right)\leq H_{1/2}(\rho_{1}\|\rho_{2})$, where we used $\ln(1-x)\leq 2x$ for all $x\in[0,\frac{1}{2}]$.
-Applying this to the lower bound of Theorem˜2, and using [7, (4.12)] gives n^∗_ε(δ,p,ρ_1,ρ_2) ≥(eε+1)lnp(1-p)δ2(eε-1)H1/2(ρ1ρ2).
-Additionally, [7, (4.25)] yields that n^∗_ε(δ,p,ρ_1,ρ_2) ≥(1-δ/p)2e-ε(eε-1)2E1(ρ1ρ2)2.
+Using [8, Lemma 1], we have that $\ln F(\rho_{1},\rho_{2})\leq-\ln\left(1-\frac{1}{2}H_{1/2}(\rho_{1}\|\rho_{2})\right)\leq H_{1/2}(\rho_{1}\|\rho_{2})$, where we used $\ln(1-x)\leq 2x$ for all $x\in[0,\frac{1}{2}]$.
+Applying this to the lower bound of Theorem˜2, and using [8, (4.12)] gives n^∗_ε(δ,p,ρ_1,ρ_2) ≥(eε+1)lnp(1-p)δ2(eε-1)H1/2(ρ1ρ2).
+Additionally, [8, (4.25)] yields that n^∗_ε(δ,p,ρ_1,ρ_2) ≥(1-δ/p)2e-ε(eε-1)2E1(ρ1ρ2)2.
 The proof then follows from noting that $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$ for any $(\rho_{1},\rho_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$, whereby we get $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq\sup_{\rho_{i}\in\mathcal{D}_{i}}n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$.
 ∎
-The above result directly generalizes the sample complexity of locally differentially private simple QHT, recovering [7, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
+The above result directly generalizes the sample complexity of locally differentially private simple QHT, recovering [8, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
 However, as in the previous section, the real challenge is to derive tight upper bounds on the sample complexity. To this end, we derive the following result.
@@
 Proof:
-Let $\mathcal{B}$ be the $\varepsilon$-LDP channel as constructed in the proof of [7, Theorem 4.2]. Then, we have
+Let $\mathcal{B}$ be the $\varepsilon$-LDP channel as constructed in the proof of [8, Theorem 4.2]. Then, we have
 $-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}\ln F^{\mathcal{M}}_{\max}$ $\geq 1-\inf_{\begin{subarray}{c}\mathcal{M}\in\operatorname{LDP}_{\varepsilon}\end{subarray}}F^{\mathcal{M}}_{\max}$
@@
 $\geq\frac{1}{2}\left(\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)^{2}\inf_{\rho_{i}\in\mathcal{D}_{i}}E_{1}(\rho_{1}\|\rho_{2})^{2},$
-where the first inequality uses $1-x\leq-\ln x$ for all $x\in[0,1]$, the second uses [7, Equation 4.7], and the last inequality uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$.
+where the first inequality uses $1-x\leq-\ln x$ for all $x\in[0,1]$, the second uses [8, Equation 4.7], and the last inequality uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$.
 ∎